OpenStudy (anonymous):

What is the factored form of x^2+6x+8? What is the factored form of x^2-7x+12 The area of a garden is given by the trinomial g^2-2g-24. The garden's length is g+4. What is the garden's width?

OpenStudy (triciaal):

What are the factors of 8 such that when you add them you get 6?

OpenStudy (anonymous):

2 right?

OpenStudy (triciaal):

what are the factors of 8

OpenStudy (triciaal):

8,1 and 4.2 you want the sum to be 6 so we will use 4,2

OpenStudy (triciaal):

(x +4)(x +2)

OpenStudy (triciaal):

x^2 -7x+12 what are the factors of 12 that add up to 7? x(x-4)-3(x-4) (x-3)(x-4)

OpenStudy (anonymous):

That's for the first ? right?

OpenStudy (triciaal):

I did #1 and #2

OpenStudy (anonymous):

Thank u so much! :)

OpenStudy (triciaal):

are you ready for #3?

OpenStudy (triciaal):

any ?

OpenStudy (triciaal):

the area is the length x the width A = LW = g^2-2g-24. length ,L = g + 4 W = A/L = (g^2-2g-24)./g + 4 hint: g + 4 is a factor of the expression for the area

OpenStudy (anonymous):

Thank you (: